OpenStudy (anonymous):

What is the length of the major axis? (x-7)^2/4 + (y+3)^2/16=1

OpenStudy (jdoe0001):

\(\bf \cfrac{(x-7)^2}{4}+\cfrac{(y+3)^2}{16}=1\implies \cfrac{(x-7)^2}{2^2}+\cfrac{(y+3)^2}{4^2}=1\\ \quad \\ \implies \cfrac{(x-h)^2}{a^2}+\cfrac{(y-k)^2}{b^2}=1\) 

OpenStudy (jdoe0001):

hmmm is really a vertical ellipse.... but anyhow, that's how you'd get the major axis length either way

OpenStudy (jdoe0001):